How many moles are in 8.33 x 1024 molecules of Carbon?
1
Expert's answer
2020-11-04T14:23:27-0500

A mole of carbon molecules contains 6.02×1023 molecules.

Proportion:

1 mol - 6.02×1023 molecules

x mol - 8.33 × 1024 molecules

"x = \frac{8.33 \times 10^{24}}{6.02 \times 10^{23}}= 13.83 \;mol"

Answer: 13.83 mol